# Non-paraxial effects on laser-qubit interactions

### v1

@@ III Simulated light potential
Since the potential is primarily due to the Rabi frequency between the qubit states, it follows the same spatial profile. For $B_{x}$ the potential is suppressed and spatially-symmetric about the origin. For $B_{y}$ we note that while the profile is Gaussian, it is shifted from the atom’s center-of-mass position. Moreover, we plot the potentials for a Laguerre-Gauss mode $\text{LG}_{p,l}=\text{LG}_{01}$ in Fig. 3(e)-(h), and observe a similar shifted potential in the $B_{y}$ case. We note that for the set of parameters considered here, the Rabi frequency shifts $\Omega_{s}$ due to off-resonant Raman coupling are negligible and of the order of $100$ Hz. The light shifts due to off-resonant dipole coupling are also small, on the order of 1 kHz, and their effect on the potential profile can also be neglected.
The displaced potentials observed in Fig. 3(b) & (f) are due to circularly-polarized components, arising from the strong curvature of the longitudinal field near the focus of the tweezer, which are not described in the paraxial approximation. We have previously studied the emergence of the optical Magnus effect in dipole transitions where an off-axis field generates a transverse force on the atom [22]. Here we describe a similar phenomenon when driving quadrupole transitions.
@@
We explore the dependence of the size of the Rabi frequency displacement on the tweezer waist. To this end, we determine the position of the peak ($x_{0}$) of $\Omega(x)$ by solving $\frac{\partial\Omega}{\partial x}|_{x=x_{0}}=0$ to characterize how $x_{0}$ shifts in the direction of the focal plane. We consider the case of (i) $\mathbf{B}=B_{y}\parallel\varepsilon_{y}$ and (ii) $\mathbf{B}=B_{x}\perp\varepsilon_{y}$. In the first case, the coupling $\Omega_{E2,1/2,3/2}(x)$ is resonant and $x_{0}$ approaches $k=\lambda/2\pi$ as $w_{0}\gg\lambda$, as shown in Fig. 4. In the second case, the displacement of resonant coupling $\Omega_{E2,1/2,5/2}(x)$ approaches $2k=\lambda/\pi$ as $w_{0}\gg\lambda$ (see Fig. 4). In both cases, for very small beam waists $w_{0}\ll\lambda$ (which are typically below the diffraction limit) $x_{0}$ decreases. Finally, we note when $q$ is negative the peak is displaced to the left instead of to the right (i.e. for $\Omega_{E2,1/2,-1/2}$, $x_{0}$ approaches $-\lambda/2\pi$ as $w_{0}\gg\lambda$).
The AC Stark shifts $\delta_{i}$ are also shown in Fig. 3. We only show the $B_{y}$ case due to the Stark shifts being much smaller for $B_{x}$. We note that the spatial profile is different from the Rabi frequency and in particular, the displacement is in the opposite direction. This is due to the dominant contribution to $\delta_{i}$ ($\Omega_{E2,1/2,-1/2}$) being centered at $x\approx-\lambda/2\pi$. The Stark shift scales with the square of the Rabi frequency amplitude, $\delta\propto|\Omega|^{2}$, and is inversely proportional to the magnetic field strength $|\mathbb{B}|$. For an atom positioned at the peak of the Rabi frequency, this displacement would lead to a Stark shift gradient at the atom’s position - a situation we explore in section IV.
@@
Figure 4: Displacement $x_{0}$ of the peak of the Rabi frequency as a function of waist size $w_{0}$. The scale is logarithmic on the horizontal axis. When $\mathbb{B}\parallel\varepsilon_{y}$, the resonant coupling $\Omega_{E2,1/2,3/2}$ is displaced by $x_{0}=\lambda/2\pi$ (blue line) in the limit $w_{0}\gg\lambda$. When $\mathbb{B}\perp\varepsilon_{y}$ the resonant coupling $\Omega_{E2,1/2,5/2}$ is displaced by $x_{0}=\lambda/\pi$ (orange line) in the same limit.
IV Effects on qubit manipulation
IV.1 Analytic model
The spatially varying Rabi frequency and Stark shift may be a source of error in single qubit gates. We investigate their effect by expanding both up to second order in position, $\delta(\hat{x})=\delta^{(0)}+\delta^{(1)}\hat{x}+\delta^{(2)}\hat{x}^{2}$ and $\Omega(\hat{x})=\Omega^{(0)}+\Omega^{(1)}\hat{x}+\Omega^{(2)}\hat{x}^{2}$. Here we ignore a global Stark shift imposing $\delta_{1}(\hat{x})=-\delta_{2}(\hat{x})$. We place the ion at $x_{0}$ which sets the Rabi frequency gradient $\Omega^{(1)}=0$. The Hamiltonian describing the system is then,
where $\hat{x}=l_{\text{ho}}(\hat{a}+\hat{a}^{\dagger})$ and $\hat{p}=i(\hbar/2l_{\text{ho}})(\hat{a}^{\dagger}-\hat{a})$, $\hat{a}\,(\hat{a}^{\dagger})$ is the annihilation (creation) operator with $l_{\text{ho}}=\sqrt{\hbar/2m\omega}$ the characteristic length of the harmonic oscillator. The errors associated with qubit-motion coupling due to Rabi frequency curvature $\Omega^{(2)}$ have been studied in previous works [29, 42]. We note that errors due to the Stark shift gradient $\delta^{(1)}$ only arise in the non-paraxial case, when the peak of the Stark shift lies off-axis.
We apply two unitary transformations on $H_{3}$ (see Appendix A) to eliminate the linear and quadratic qubit-motion coupling induced by the Stark shift components $\delta^{(1)}$ and $\delta^{(2)}$ respectively. Here we neglect off-resonant motional processes that change the Fock state. The transformed Hamiltonian $\tilde{H}$ in angular frequency units is,
where $H_{\text{id}}=\omega(\hat{a}^{\dagger}a+1/2)+\Omega^{(0)}\hat{\sigma}_{x}$ is an ideal $\hat{\sigma}_{x}$ qubit rotation. The second term can be interpreted as the error from squeezing the ion’s harmonic motion due to the Stark shift curvature, labeled with error parameter $\kappa_{z}=2\delta^{(2)}l_{\text{ho}}^{2}$. The third term describes $\sigma_{x}$-type errors due to the Rabi frequency curvature and the Stark shift gradient, labeled $\kappa_{x}=2\Omega^{(0)}(\delta^{(1)})^{2}l_{\text{ho}}^{2}/\omega^{2}+\Omega^{(2)}l_{\text{ho}}^{2}$. The fidelity of the resulting gate is given by [43],
$\bar{F}=\frac{\sum_{l}tr[\hat{U}_{\text{id}}\hat{\sigma}^{\dagger}_{l}\hat{U}_{\text{id}}^{\dagger}\boldsymbol{\hat{\sigma}}_{l}(\hat{U}_{\text{real}})+d^{2}]}{d^{2}(d+1)},$ (18)
where $\boldsymbol{\hat{\sigma}}_{l}(\hat{U}_{\text{real}})\equiv\text{tr}_{\text{Fs}}\left(\hat{U}_{\text{real}}(\ket{n}\bra{n}\otimes\hat{\sigma}_{l})\hat{U}_{\text{real}}^{\dagger}\right)$ is the projector on Fock state |$\ket{n}$, and $d=2$ [16]. The unitaries are $\hat{U}_{\text{id}}=\exp(-iH_{\text{id}}t)$ and $\hat{U}_{\text{real}}=\exp(-i\tilde{H}t)$, and the motional modes are weighted with the Bose-Einstein distribution,
$P_{\text{th}}(n,\bar{n})=\frac{\bar{n}^{n}}{(1+\bar{n})^{n+1}}.$ (19)
Here the average occupation number $\bar{n}=1/(e^{\frac{\omega\hbar}{k_{B}T}}-1)$ is determined by the temperature of the ion crystal, $T$.
@@ IV.1 Analytic model
Considering the physical contributions of the infidelity, the leading error parameter $\kappa_{x}$ is dominated by the Rabi frequency curvature $\Omega^{(2)}$ when $\Omega^{(2)}/\Omega^{(0)}>2(\delta^{(1)})^{2}/\omega^{2}$. The ratio of the Rabi frequency curvature to amplitude $\Omega^{(2)}/\Omega^{(0)}$ only changes geometrically, for instance when the waist is decreased, which increases the Stark shifts quadratically. Consequently, when the trap frequency $\omega$ or magnetic field strength is lower, or the beam intensity is higher, the non-paraxial error becomes more significant. For instance, when $\Omega^{(0)}=2\pi~{}1$ MHz, $\mathbb{B}=2.5$ G and $\omega_{z}=2\pi~{}40$ kHz, the non-paraxial $\delta^{(1)}$ error surpasses the $\Omega^{(2)}$ error, although these parameters are unusual for trapped ion experiments.
IV.2 Compensated gate time
@@
The $\kappa_{x}$ error is an error in the timing of the qubit rotation and can be partially compensated. Assuming we know $\bar{n}$ we can define a new gate time $\bar{t}_{g}=\pi/2\bar{\Omega}$, where
@@
where $\gamma=\Omega^{(0)}/\bar{\Omega}$. The analytic expression above allow us to significantly reduce the infidelity without numerical optimization of the gate time. The fidelity calculated with both gate times is compared against numerical simulation in Fig. 5.
Figure 5: Simulated infidelity for a $\pi/2$ gate as a function of average occupation $\bar{n}$. We compare simulated errors due to the Rabi frequency $\Omega(x)$ (blue points) and additionally the Stark shift $\delta(x)$ (orange points), with the ion placed at the peak of the Rabi frequency $x_{0}$ and using a gate time $\bar{t}_{g}$. The dashed lines are analytic expressions for the infidelity with a compensated (purple dashed) and uncompensated (red dotted) gate time. We consider Rabi frequency amplitudes $\Omega^{(0)}=2\pi~{}0.2$ MHz and $12\pi~{}$ 1 MHz, and magnetic field strengths $|\textbf{B}|=2.5$ G and $5$ G. Decreasing the magnetic field linearly increases the Stark shift error, making it more prominent. Increasing the Rabi frequency increases off-resonant errors, which dominate at low $\bar{n}$.
IV.3 Numerical simulation
We compute the fidelity numerically with the Hamiltonian in Eq. (16). In Figure 5, we show the infidelity evaluated for the parameters in the previous section, with an axial trap frequency $\omega_{z}=2\pi~{}0.5$ MHz corresponding to $l_{\text{ho}}=16$ nm, with two Rabi frequency amplitudes $\Omega^{(0)}=2\pi~{}0.2$ MHz and $12\pi~{}$ 1 MHz, and two magnetic field strengths $|\mathbb{B}|=2.5$ G and $5$ G. For the case $\Omega^{(0)}=2\pi~{}1$ MHz and $|\mathbb{B}|=5$ G, the field gradients and curvatures at the position of the ion (at $x_{0}\sim 108$ nm from the beam center) are $\delta^{(1)}\approx 2\pi~{}20$ Hz/nm, $\delta^{(2)}\approx 2\pi~{}0.03$ Hz/nm2 and $\Omega^{(2)}\approx 2\pi~{}2$ Hz/nm2, corresponding to fractional error parameters $\kappa_{x}/\Omega^{(0)}\approx 5\times 10^{-4}$ and $\kappa_{z}/\Omega^{(0)}\approx 1\times 10^{-5}$. We find that $\Omega^{(2)}$ is the dominant error source, as evidenced by the small difference between the solid orange and blue lines, which include and neglect $\delta(x)$, respectively.
V Conclusions
In this paper we considered the effects of light potentials from an optical tweezer on a trapped ion beyond the paraxial approximation. The spatial profile of the light potentials of Gaussian and Laguerre-Gaussian beams are shifted off-axis with respect to the beam intensity. The shift is determined by the orientation of the magnetic field with respect to the polarization, and the change in angular momentum $q$ for the quadrupole transition. We showed that when the ion is placed at the peak of the Rabi frequency $\Omega(x)$, it experiences a non-zero curvature $\Omega(x)$, and a non-zero gradient of the Stark shift $\delta(x)$.
Next, we evaluated the size of the errors resulting from the non-zero curvature and gradient terms, which manifest as qubit-motion entanglement. Our numerical simulations, as well as analytic results, show that while the size of the error is strongly dependent on the trapping frequency and the ion temperature, the non-paraxial errors are generally very small when using typical experimental parameters. In particular, the errors arising from the effects considered here lie below the typical thresholds for quantum error correction [44], and thus do not pose a serious threat to manipulating ion qubits with tightly-focused beams. However, at low temperatures and high laser intensities non-paraxial effects can become significant. This may be important for the next-generation of ion trap quantum computers, which strive towards lower temperatures and utilize individually-addressing, high-intensity beams. Furthermore, our theoretical treatment can be applied towards future work, for instance studying non-paraxial effects on neutral atoms, and for engineering quantum logic gates with optical tweezers.
@@ Acknowledgements
We would like to acknowledge Liam J. Bond for help on the Hamiltonian transformations. This work was supported by the Netherlands Organization for Scientific Research (Grant Nos. 680.91.120, VI.C.202.051 and 680.92.18.05), the Dutch Research Council (Grant No. OCENW.M.22.403), and by the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via project 101070144 (EuRyQa) A.S.N. is supported by the Dutch Research Council (NWO/OCW) as part of the Quantum Software Consortium programme (project number 024.003.037). A.S.N. is supported by Quantum Delta NL (project number NGF.1582.22.030).
Appendix A Squeeze and displacement transformation
The Hamiltonian of the system is
$H_{3}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega^{2}\hat{x}^{2}+\hbar\left(\delta^{(0)}+\delta^{(1)}\hat{x}+\delta^{(2)}\hat{x}^{2}\right)\sigma_{z}+\hbar\left(\Omega^{(0)}+\Omega^{(2)}\hat{x}^{2}\right)\sigma_{x}$ (23)
where the $\hat{x}$ and $\hat{p}$ operators are defined in terms of the creation and annihilation operators $\hat{x}=l_{\text{ho}}(\hat{a}+\hat{a}^{\dagger})$ and $\hat{p}=i(\hbar/2l_{\text{ho}})(\hat{a}^{\dagger}-\hat{a})$, and $l_{\text{ho}}=\sqrt{\hbar/2m\omega}$.
@@ Appendix A Squeeze and displacement transformation
$\hat{S}(\xi\hat{\sigma}_{z})=\exp(\frac{\xi\hat{\sigma}_{z}}{2}(\hat{a}^{2}-\hat{a}^{\dagger}{}^{2})),$ (24)
Using the quadrature transformations $\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{a}\hat{S}(\xi\hat{\sigma}_{z})=\hat{a}\cosh(\xi\hat{\sigma}_{z})-\hat{a}^{\dagger}\sinh(\xi\hat{\sigma}_{z})$ and $\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{a}^{\dagger}\hat{S}(\xi\hat{\sigma}_{z})=\hat{a}^{\dagger}\cosh(\xi\hat{\sigma}_{z})-\hat{a}\sinh(\xi\hat{\sigma}_{z})$, the motional terms transform as
$\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{x}^{2}\hat{S}\xi\hat{\sigma}_{z})$ $=e^{-2\xi\hat{\sigma}_{z}}\hat{x}^{2},$ (25a)
$\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{x}\hat{S}(\xi\hat{\sigma}_{z})$ $=e^{-\xi\hat{\sigma}_{z}}\hat{x},$ (25b)
$\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{p}^{2}\hat{S}(\xi\hat{\sigma}_{z})$ $=e^{2\xi\hat{\sigma}_{z}}\hat{p}^{2}.$ (25c)
@@
$\tilde{H}$ $=\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})H_{3}\hat{S}(\xi\hat{\sigma}_{z})$ (26)
$=\frac{1}{2m}e^{2\xi\hat{\sigma}_{z}}\hat{p}^{2}+\frac{1}{2}e^{-2\xi\hat{\sigma}_{z}}m\omega^{2}\hat{x}^{2}+\hbar\left(\delta^{(0)}+\delta^{(1)}e^{-\xi\hat{\sigma}_{z}}\hat{x}+\delta^{(2)}e^{-2\xi\hat{\sigma}_{z}}\hat{x}^{2}\right)\hat{\sigma}_{z}+\hbar\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z})$ (27)
@@
We define a qubit-dependent frequency $\tilde{\omega}=\sqrt{\omega^{2}+\frac{2\delta^{(2)}\hbar}{m}\hat{\sigma}_{z}}$, which absorbs the quadratic $\delta^{(2)}$ shift into the ion’s harmonic motion. By 1st-order Taylor expansion $\tilde{\omega}\approx\omega+(\delta^{(2)}l_{\text{ho}}^{2})\hat{\sigma}_{z}$. The Rabi frequency components don’t commute with the squeezing operator, so we have abbreviated them $\hat{\Omega}=\Omega^{(0)}+\Omega^{(2)}\hat{x}^{2}$ for now. Writing $\hat{x}$ and $\hat{p}$ in terms of $\hat{a}$ and $\hat{a}^{\dagger}$, the Hamiltonian (in angular frequency units) is
$\tilde{H}=$ $-\frac{e^{2\xi\hat{\sigma}_{z}}}{4}\omega\left(\hat{a}^{\dagger^{2}}+\hat{a}^{2}-2\hat{a}^{\dagger}\hat{a}-1\right)+\frac{e^{-2\xi\hat{\sigma}_{z}}}{4}\frac{\tilde{\omega}^{2}}{\omega}\left(\hat{a}^{\dagger^{2}}+\hat{a}^{2}+2\hat{a}^{\dagger}\hat{a}+1\right)$
$+\left(\delta^{(0)}+\delta^{(1)}e^{-\xi\hat{\sigma}_{z}}l_{\text{ho}}(\hat{a}+\hat{a}^{\dagger})\right)\hat{\sigma}_{z}+\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z})$ (28)
@@
We recover the harmonic oscillator with frequency $\tilde{\omega}$ by choosing the squeezing parameter $\xi=\delta^{(2)}l_{\text{ho}}^{2}/\omega$, making it a dimensionless measure of the Stark shift curvature. The Hamiltonian is then
$\tilde{H}=\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)+\delta^{(0)}\hat{\sigma}_{z}+\delta^{(1)}e^{-{\xi\hat{\sigma}_{z}}}x_{0}(\hat{a}+\hat{a}^{\dagger})\hat{\sigma}_{z}+\hat{S}^{\dagger}(\xi\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(\xi\hat{\sigma}_{z}).$ (29)
@@
Similarly the linear motional term can be transformed through a qubit-dependent Lang-Firsov transformation with unitary
$\hat{D}(\alpha\hat{\sigma}_{z})=\exp(\alpha\hat{\sigma}_{z}(\hat{a}^{\dagger}-\hat{a})),$ (30)

### v2

@@ IV.2 Simulated light potentials
In Fig.  5 we determine the position of the peak ($x_{0}$) of $\Omega(x)$ to characterize how $x_{0}$ shifts in the direction of the focal plane. We consider the case of (i) $\mathbf{B}=B_{y}\parallel\varepsilon_{y}$ and (ii) $\mathbf{B}=B_{x}\perp\varepsilon_{y}$. In the first case, the coupling $\Omega_{E2,1/2,3/2}(x)$ is resonant and $x_{0}$ approaches $k^{-1}=\lambda/2\pi$ as $w_{0}\gg\lambda$, as shown in Fig. 5. In the second case, the displacement of resonant coupling $\Omega_{E2,1/2,5/2}(x)$ approaches $2k^{-1}=\lambda/\pi$ as $w_{0}\gg\lambda$ (see Fig. 5). In both cases, for very small beam waists $w_{0}\ll\lambda$ (which are typically below the diffraction limit) $x_{0}$ decreases. Finally, we note when $q$ is negative the peak is displaced to the left instead of to the right (i.e. for $\Omega_{E2,1/2,-1/2}$, $x_{0}$ approaches $-\lambda/2\pi$ as $w_{0}\gg\lambda$).
@@
Figure 5: Rabi frequency displacement. Displacement $x_{0}$ of the peak of the Rabi frequency $\Omega(x)$ as a function of waist size $w_{0}$. The scale is logarithmic on the horizontal axis. When $\mathbb{B}\parallel\varepsilon_{y}$, the resonant coupling $\Omega_{1/2\rightarrow 3/2}(x)$ is displaced by $x_{0}=\lambda/2\pi$ (blue line) in the limit $w_{0}\gg\lambda$. When $\mathbb{B}\perp\varepsilon_{y}$ the resonant coupling $\Omega_{1/2\rightarrow 5/2}(x)$ is displaced by $x_{0}=\lambda/\pi$ (orange line) in the same limit.
We plot the the spatial profile of the AC Stark shifts $\delta_{i}$ for $\mathbf{B}=B_{y}\hat{y}$ in 2. The Stark shifts for $\mathbf{B}=B_{x}\hat{x}$ are negligible in comparison. We note that the spatial profile is different from the Rabi frequency and in particular, the displacement is in the opposite direction. This is due to the dominant contribution to $\delta_{i}$ ($\Omega_{E2,1/2,-1/2}$) being centered at $x\approx-\lambda/2\pi$. The Stark shift scales with the square of the Rabi frequency amplitude, $\delta\propto|\Omega|^{2}$, and is inversely proportional to the magnetic field strength $|\mathbb{B}|$. For an atom positioned at the peak of the Rabi frequency, this displacement results in a Stark shift gradient at the atom’s position.
IV.3 Single-qubit infidelity
@@
The largest contribution to the infidelity in Eq. (5) is $\kappa_{x}$, which is predominantly due to the Rabi frequency curvature $\Omega^{(2)}$ when $\Omega^{(2)}/\Omega^{(0)}>2(\delta^{(1)})^{2}/\omega^{2}$. We note that the ratio of the Rabi frequency curvature to its amplitude $\Omega^{(2)}/\Omega^{(0)}$ changes geometrically, for instance as the beam waist is decreased. This in turn increases the Stark shifts quadratically which means that for sufficiently small trap frequency $\omega$ or magnetic field strength, and at large enough beam intensities, non-paraxial effects are the dominant source of error. One such choice of parameters is $\Omega^{(0)}=2\pi~{}1$ MHz, $\mathbb{B}=2.5$ G and $\omega=2\pi~{}40$ kHz, where the non-paraxial $\delta^{(1)}$ error surpasses the $\Omega^{(2)}$ error. However one should keep in mind that these parameters are unusual for trapped ion experiments.
@@
We partially compensate the $\kappa_{x}$ error by using a gate time $\bar{t}_{g}=\pi/2\bar{\Omega}$, where
@@
where $\gamma=\Omega^{(0)}/\bar{\Omega}$.
IV.4 Parallel single-qubit gates
We consider addressing two ions in an $N$-ion chain simultaneously in order to perform parallel single-qubit gates. In the interaction picture with respect to $H_{0}=\omega_{m}\hat{a}^{\dagger}_{m}\hat{a}_{m}$, the Hamiltonian is:
$H_{4}=\Omega^{(0)}(\hat{\sigma}_{x}^{(i)}+\hat{\sigma}_{x}^{(j)})+\Omega^{(2)}\left(\hat{x}_{i}^{2}\hat{\sigma}_{x}^{(i)}+\hat{x}_{j}^{2}\hat{\sigma}_{x}^{(j)}\right)$
$+\delta^{(1)}\left(\hat{x}_{i}\hat{\sigma}_{z}^{(i)}+\hat{x}_{j}\hat{\sigma}_{z}^{(j)}\right)+\delta^{(2)}\left(\hat{x}_{i}^{2}\hat{\sigma}_{z}^{(i)}+\hat{x}_{j}^{2}\hat{\sigma}_{z}^{(j)}\right).$ (20)
The first term of Hamiltonian (20) describes parallel $\hat{\sigma}_{x}$ qubit rotations. Performing a Lang-Firsov transformation on $H_{4}$ with unitary $\hat{U}_{1}=\exp(\sum_{m}\zeta_{m}(b_{mi}\hat{\sigma}_{z}^{(i)}+b_{mj}\hat{\sigma}_{z}^{(j)})(\hat{a}^{\dagger}_{m}-\hat{a}_{m}))$ where $\zeta_{m}=\delta^{(1)}l_{m}/\omega_{m}$ eliminates the linear phonon coupling. Disregarding global energy offsets and off-resonant processes that change the Fock state, the transformed Hamiltonian is
$H_{5}=$ $\Omega^{(0)}(1+2\zeta_{m}^{2}(2\hat{n}_{m}+1))(b_{mi}\hat{\sigma}_{x}^{(i)}+b_{mj}\hat{\sigma}_{x}^{(j)})$
$+\Omega^{(2)}l_{m}^{2}(2\hat{n}_{m}+1)(b_{mi}\hat{\sigma}_{x}^{(i)}+b_{mj}\hat{\sigma}_{x}^{(j)})$
$+\delta^{(2)}l_{m}^{2}(2\hat{n}_{m}+1)(b_{mi}\hat{\sigma}_{z}^{(i)}+b_{mj}\hat{\sigma}_{z}^{(j)})$
$-2\omega_{m}\zeta_{m}^{2}b_{bi}b_{mj}\hat{\sigma}_{z}^{(i)}\hat{\sigma}_{z}^{(j)}.$ (21)
Here we have used the Baker-Haussdorf lemma to determine the leading-order corrections as in the single qubit case (see the Supplemental Material). This Hamiltonian describes qubit-motion coupling of the same form as the single-qubit case, albeit now coupling to each axial mode. We note an additional qubit-qubit interaction (final term), making operations on qubit $i$ dependent on the state of qubit $j$.
@@ Acknowledgements
We thank Liam J. Bond for help on the Hamiltonian transformations. This work was supported by the Netherlands Organization for Scientific Research (Grant Nos. 680.91.120, VI.C.202.051 and 680.92.18.05), the Dutch Research Council (Grant No. OCENW.M.22.403), and by the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via project 101070144 (EuRyQa) A.S.N. is supported by the Dutch Research Council (NWO/OCW) as part of the Quantum Software Consortium programme (project number 024.003.037). A.S.N. is supported by Quantum Delta NL (project number NGF.1582.22.030).
V Author contributions
LG performed the analytic calculations and wrote the manuscript with support from RS. LG and MM performed the numerical simulation of gate errors. ZA supported the numerical simulations and interpretation of results. RG, ASN and RS conceptually developed and supervised the project. All authors contributed to the manuscript.
S1 Gradient and curvature terms
Here we give analytic expressions for the gradient and curvature terms in $H_{3}$ which cause the considered errors in the main text. The Rabi frequency $\Omega(x)$ is approximately a displaced Gaussian, and we make a Taylor expansion about the Rabi frequency peak $x_{0}$ (where the ion is positioned):
$\Omega(\hat{x})$ $\approx\Omega^{(0)}\exp\left(-\frac{(\hat{x}-x_{0})^{2}}{w_{0}^{2}}\right)$
$\approx\Omega^{(0)}\left(1-\frac{(\hat{x}-x_{0})^{2}}{w_{0}^{2}}+\dots\right).$ (S22)
We identify the coefficient of $\hat{x}^{2}$ as: $\Omega^{(2)}\approx\frac{\Omega^{(0)}}{w_{0}^{2}}$. Similarly, the Stark shift is a displaced Gaussian but now shifted in the opposite direction and centered at $-x_{0}$. The Taylor expansion about the ion’s position yields:
$\delta(\hat{x})$ $\approx\delta^{(0)}\exp\left(-\frac{(\hat{x}+x_{0})^{2}}{w_{0}^{2}}\right)$
$\approx\delta^{(0)}\exp(-4x_{0}^{2}/w_{0}^{2})\biggl{(}1-\frac{4x_{0}}{w_{0}^{2}}(\hat{x}-x_{0})~{}$
$~{}~{}-\frac{(w_{0}^{2}-8x_{0}^{2})}{w_{0}^{4}}(\hat{x}-x_{0})^{2}+\dots\biggr{)}.$ (S23)
For larger waists with $w_{0}^{2}\gg x_{0}^{2}$ such that $\exp(-4x_{0}^{2}/w_{0}^{2})\sim 1$, the gradient and curvature terms simplify to $\delta^{(1)}\sim\frac{4\delta^{(0)}x_{0}}{w_{0}^{2}}$ and $\delta^{(2)}\sim\frac{\delta^{(0)}(w_{0}^{2}-8x_{0}^{2})}{w_{0}^{4}}$. Note that in our case with $w_{0}=\lambda$ and $x_{0}=\lambda/2\pi$, $\exp(-4x_{0}^{2}/w_{0}^{2})=\exp(-\pi^{-2})\sim 0.9$.
S2 Hamiltonian transformations
@@ V Author contributions
$\hat{S}(s\hat{\sigma}_{z})=\exp(\frac{s\hat{\sigma}_{z}}{2}(\hat{a}^{2}-\hat{a}^{\dagger}{}^{2})),$ (S24)
Using the quadrature transformations $\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{a}\hat{S}(s\hat{\sigma}_{z})=\hat{a}\cosh(s\hat{\sigma}_{z})-\hat{a}^{\dagger}\sinh(s\hat{\sigma}_{z})$ and $\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{a}^{\dagger}\hat{S}(s\hat{\sigma}_{z})=\hat{a}^{\dagger}\cosh(s\hat{\sigma}_{z})-\hat{a}\sinh(s\hat{\sigma}_{z})$, the motional terms transform as
$\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{x}^{2}\hat{S}s\hat{\sigma}_{z})$ $=e^{-2s\hat{\sigma}_{z}}\hat{x}^{2},$ (S25a)
$\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{x}\hat{S}(s\hat{\sigma}_{z})$ $=e^{-s\hat{\sigma}_{z}}\hat{x},$ (S25b)
$\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{p}^{2}\hat{S}(s\hat{\sigma}_{z})$ $=e^{2s\hat{\sigma}_{z}}\hat{p}^{2}.$ (S25c)
@@
$\tilde{H}$ $=\hat{S}^{\dagger}(s\hat{\sigma}_{z})H_{3}\hat{S}(s\hat{\sigma}_{z})$
$=\frac{1}{2m}e^{2s\hat{\sigma}_{z}}\hat{p}^{2}+\frac{1}{2}e^{-2s\hat{\sigma}_{z}}m\omega^{2}\hat{x}^{2}$
$+\hbar\left(\delta^{(0)}+\delta^{(1)}e^{-s\hat{\sigma}_{z}}\hat{x}+\delta^{(2)}e^{-2s\hat{\sigma}_{z}}\hat{x}^{2}\right)\hat{\sigma}_{z}$
$+\hbar\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z})$ (S26)
@@
We define a qubit-dependent frequency $\tilde{\omega}=\sqrt{\omega^{2}+\frac{2\delta^{(2)}\hbar}{m}\hat{\sigma}_{z}}$, which absorbs $\delta^{(2)}$ into the ion’s harmonic motion. By 1st-order Taylor expansion $\tilde{\omega}\approx\omega+(\delta^{(2)}l_{\text{ho}}^{2})\hat{\sigma}_{z}$. The Rabi frequency components are denoted $\hat{\Omega}=\Omega^{(0)}+\Omega^{(2)}\hat{x}^{2}$. In terms of creation and annihilation operators, the Hamiltonian (in angular frequency units) is
$\tilde{H}=$ $-\frac{e^{2s\hat{\sigma}_{z}}}{4}\omega\left(\hat{a}^{\dagger^{2}}+\hat{a}^{2}-2\hat{a}^{\dagger}\hat{a}-1\right)$
$+\frac{e^{-2s\hat{\sigma}_{z}}}{4}\frac{\tilde{\omega}^{2}}{\omega}\left(\hat{a}^{\dagger^{2}}+\hat{a}^{2}+2\hat{a}^{\dagger}\hat{a}+1\right)$
$+\left(\delta^{(0)}+\delta^{(1)}e^{-s\hat{\sigma}_{z}}l_{\text{ho}}(\hat{a}+\hat{a}^{\dagger})\right)\hat{\sigma}_{z}$
$+\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z})$ (S27)
@@
We recover the harmonic oscillator with frequency $\tilde{\omega}$ by choosing the squeezing parameter $s=\delta^{(2)}l_{\text{ho}}^{2}/\omega$ as a dimensionless measure of the Stark shift curvature. The Hamiltonian is then
$\tilde{H}=$ $\tilde{\omega}(\hat{a}^{\dagger}\hat{a}+1/2)+\delta^{(0)}\hat{\sigma}_{z}$
$+\delta^{(1)}e^{-{s\hat{\sigma}_{z}}}x_{0}(\hat{a}+\hat{a}^{\dagger})\hat{\sigma}_{z}$
$+\hat{S}^{\dagger}(s\hat{\sigma}_{z})\hat{\Omega}\hat{\sigma}_{x}\hat{S}(s\hat{\sigma}_{z}).$ (S28)
@@
Similarly the linear motion term can be transformed through a qubit-dependent Lang-Firsov transformation with unitary
$\hat{D}(\zeta\hat{\sigma}_{z})=\exp(\zeta\hat{\sigma}_{z}(\hat{a}^{\dagger}-\hat{a})),$ (S29)